# Accurate and Efficient Low-Rank Model Merging in Core Space

### v3

@@ 4.1 Model Merging in Core Space
$[B^{(1)};\cdots;B^{(T)}]=U_{B}^{\mathrm{ref}}\Sigma_{B}V_{B}^{\top};\quad\begin{bmatrix}A^{(1)}\\ \vdots\\ A^{(T)}\end{bmatrix}=U_{A}\Sigma_{A}\left(V_{A}^{\mathrm{ref}}\right)^{\top}.$ (3)
@@
  • Efficiency. Core matrices $M^{(t)}\in\mathbb{R}^{Tr\times Tr}$ are significantly smaller than their full-space counterparts $\Delta W^{(t)}\in\mathbb{R}^{m\times n}$. This reduction enables high-cost merging algorithms to run at a fraction of the time and memory footprint (see Sec.˜4.3).
  • Efficacy. As shown in Sec.˜5.1, Core Space merging improves performance over full-space merging when non-linear methods are used. In Sec.˜5.2, we show that this improvement stems from better alignment and more compact representation of task-specific directions.
@@ 5 Experimental Results
Experimental setup. We follow the experimental setup of KnOTS and use the LoRA checkpoints provided by the authors (stoica2024knots, ). For the vision experiments, we use two variants of CLIP (radford2021learning, ) with ViT-B/32 and ViT-L/14 dosovitskiy2021an as vision encoders fine-tuned on a standard set of 8 tasks. We employ Llama 3 8B (grattafiori2024llama, ) fine-tuned on 6 NLI tasks for the language experiments. All models are fine-tuned with LoRA (hu2021lora, ) with rank 16 applied on all matrices (keys, queries, values, and outputs) across all attention layers. Following stoica2024knots , we report normalized accuracy as a ratio of the accuracy of the merged model on a given task to the accuracy of the model fine-tuned on this task. We also report absolute accuracy for the joint-task setting (additional experimental details in Appendix˜D).
@@
Baseline merging methods. We evaluate each merging space using the following merging methods. Task Arithmetic (TA) (ilharco2023task, ) performs a scaled summation of each task matrix $W_{\text{merged}}=W_{0}+\alpha\sum_{i=1}^{T}\Delta W_{i}$. As this is a linear operation, the results of merging in each space are the same (see Eq.˜11 for Core and stoica2024knots for KnOTS). TIES (yadav2023tiesmerging, ) trims low-magnitude parameters and averages parameters with dominating sign, while DARE (yu2024language, ) preprocesses task vectors by randomly dropping a fraction of parameters and rescaling the remaining ones. TSV (tsv, ) concatenates low-rank approximations of task matrices and orthogonalizes them across tasks. CART choi2024revisiting calculates centered task vectors as a difference of fine-tuned weights from the average of all fine-tuned weights and performs task arithmetic on the low-rank approximation of these centered task vectors. Iso-C (marczak2025task, ) flattens the spectrum of singular values for a model merged with task arithmetic. As the spectrum flattening can be performed on weights merged with any merging technique, we combine Iso with other merging techniques, denoting it with +Iso-C.
@@ Appendix B Computational Complexity Analysis
Iso-C Complexity. To assess the time complexity of our approach, we begin by analyzing that of Iso-C [28], to establish a baseline for comparison. For each layer, the Iso-C procedure can be broken down into the following steps:
@@
  • SVD computation: Computing the decomposition $\Delta_{\operatorname{TA}}=U\Sigma V^{\top}$ for an $m\times n$ matrix has a complexity of $\mathcal{O}(m^{2}\cdot n+n^{3})$ [40].
@@
KnOTS Complexity. Secondly, we analyze the computational cost of the KnOTS method [42]:
@@
TSV Complexity. Then we analyze the cost of TSV [12]:
@@ C.1 High subspace alignment leads to lower interference
In this Section, we experimentally show that merging in Core Space reduces interference when merging models. We follow [53, 28] and measure the interference as the L1 distance between the final embeddings of task-specific models and the merged one. We compare the interference when merging with TSV + Iso-C in Full Space versus Core Space. For each dataset, we collect the activations from the final layer (i.e., the projection to a common vision-language space) of both the task-specific model and the merged model. We present the average distance across all the samples in the test set. We observe lower interference when merging in Core Space, highlighting its effectiveness. Note that Full Space merging, which causes higher interference, also exhibits higher SAR in Sec.˜5.2.
@@ Appendix D Additional Experiment Details
  • KnOTS LoRA Checkpoints [42]: The KnOTS repository, which provides LoRA-adapted model checkpoints and training scripts, is licensed under the MIT License. This permissive license allows for reuse and modification with proper attribution.
  • Cars196 [22]: The Cars196 dataset is available for non-commercial research purposes. Specific licensing details are not explicitly provided.
  • Describable Textures Dataset (DTD) [7]: The DTD is made available to the computer vision community for research purposes. The dataset is licensed under the Creative Commons Attribution 4.0 License (CC BY 4.0).
  • EuroSAT [14]: The EuroSAT dataset is licensed under the MIT License.
  • German Traffic Sign Recognition Benchmark (GTSRB) [41]: The GTSRB dataset is licensed under the Creative Commons Zero (CC0) Public Domain Dedication.
  • MNIST [23]: The MNIST dataset is publicly available for research purposes. Specific licensing details are not explicitly provided; users are advised to consult the dataset’s source for more information.
  • NWPU-RESISC45 [5]: The NWPU-RESISC45 dataset is licensed under the Creative Commons Attribution 4.0 License (CC BY 4.0).
  • SUN397 [49]: The SUN397 dataset is available for research purposes only. Specific licensing details are not explicitly provided; users are advised to consult the dataset’s source for more information.
  • Street View House Numbers (SVHN) [33]: The SVHN dataset is available for non-commercial use only.
  • Stanford Natural Language Inference (SNLI) [3]: The SNLI dataset is licensed under the Creative Commons Attribution-ShareAlike 4.0 International License (CC BY-SA 4.0).
  • Multi-Genre Natural Language Inference (MNLI) [47]: The MNLI dataset is released under the Open American National Corpus (OANC) license, which permits free use, modification, and sharing under permissive terms.
  • Sentences Involving Compositional Knowledge (SICK) [30]: The SICK dataset is distributed under a Creative Commons Attribution-NonCommercial-ShareAlike license.
  • Question Natural Language Inference (QNLI) [45]: The QNLI dataset is part of the GLUE benchmark. Specific licensing details are not explicitly provided; users are advised to consult the dataset’s source for more information.
  • Recognizing Textual Entailment (RTE) [45]: The RTE dataset is part of the GLUE benchmark. Specific licensing details are not explicitly provided; users are advised to consult the dataset’s source for more information.
  • SciTail [20]: The SciTail dataset is licensed under the Apache License 2.0.
@@ D.1 Experimental Environment
Our implementation builds directly on the KnOTS codebase [42] and uses the exact LoRA checkpoints they released. For full details on the original training and adaptation procedures, please refer to [42].
@@ D.2 Hyperparameter Search
To find optimal hyperparameters for each model, we adopt the widely used validation holdout strategy [42, 28, 12, 50]. Specifically, we perform a linear search for hyperparameters on the validation set, starting from a defined minimum value and incrementally increasing it until performance declines, indicating the optimal range. The identified optimal hyperparameters are then applied to the test set. We use the following search settings:
@@
  • For CART, the pruning rank is searched over the set $\{0.04,0.08,0.16,0.32\}$, following the methodology of the original paper. Additionally, CART includes an extra scaling factor $\lambda$ in its merging formulation. Specifically, the merged weights are computed as $W_{\text{merged}}=W_{0}+\alpha(\theta_{\text{avg}}+\lambda\sum_{t=1}^{T}\bar{\tau}_{t})$, where $\theta_{\text{avg}}$ denotes the average of the updates and $\bar{\tau}_{t}$ represents the centered task vector for task $t$. For further details, we refer the reader to [6].
@@ E.3 Extension to VeRA
We also evaluated the applicability of Core Space merging beyond LoRA, specifically on VeRA [21] (Tab.˜10). In VeRA, the decomposition $\Delta W=\Lambda_{b}B\Lambda_{d}A$ differs structurally from LoRA since $A$ and $B$ are fixed random matrices, and only the scaling vectors $\Lambda_{b},\Lambda_{d}$ are trainable.

### v4

@@ 4.1 Model Merging in Core Space
$[B^{(1)};\ldots;B^{(T)}]=U_{B}^{\mathrm{ref}}\Sigma_{B}V_{B}^{\top};\quad\begin{bmatrix}A^{(1)}\\ \vdots\\ A^{(T)}\end{bmatrix}=U_{A}\Sigma_{A}\left(V_{A}^{\mathrm{ref}}\right)^{\top}.$ (3)
@@
  • Efficiency. Core matrices $M^{(t)}\in\mathbb{R}^{Tr\times Tr}$ are significantly smaller than their full space counterparts $\Delta W^{(t)}\in\mathbb{R}^{m\times n}$. This reduction enables high-cost merging algorithms to run at a fraction of the time and memory footprint (see Sec.˜4.3).
  • Efficacy. As shown in Sec.˜5.1, Core Space merging improves performance over full space merging when non-linear methods are used. In Sec.˜5.2, we show that this improvement stems from better alignment and more compact representation of task-specific directions.
@@ 5 Experimental Results
Experimental setup. We follow the experimental setup of KnOTS and use the LoRA checkpoints provided by the authors (stoica2024knots). For the vision experiments, we use two variants of CLIP (radford2021learning) with ViT-B/32 and ViT-L/14 dosovitskiy2021an as vision encoders fine-tuned on a standard set of 8 tasks. We employ Llama 3 8B (grattafiori2024llama) fine-tuned on 6 NLI tasks for the language experiments. All models are fine-tuned with LoRA (hu2021lora) with rank 16 applied on all matrices (keys, queries, values, and outputs) across all attention layers. Following stoica2024knots, we report normalized accuracy as a ratio of the accuracy of the merged model on a given task to the accuracy of the model fine-tuned on this task. We also report absolute accuracy for the joint-task setting (additional experimental details in Appendix˜D).
@@
Baseline merging methods. We evaluate each merging space using the following merging methods. Task Arithmetic (TA) (ilharco2023task) performs a scaled summation of each task matrix $W_{\text{merged}}=W_{0}+\alpha\sum_{i=1}^{T}\Delta W_{i}$. As this is a linear operation, the results of merging in each space are the same (see Eq.˜11 for Core and stoica2024knots for KnOTS). TIES (yadav2023tiesmerging) trims low-magnitude parameters and averages parameters with dominating sign, while DARE (yu2024language) preprocesses task vectors by randomly dropping a fraction of parameters and rescaling the remaining ones. TSV (tsv) concatenates low-rank approximations of task matrices and orthogonalizes them across tasks. CART choi2024revisiting calculates centered task vectors as a difference of fine-tuned weights from the average of all fine-tuned weights and performs task arithmetic on the low-rank approximation of these centered task vectors. Iso-C (marczak2025task) flattens the spectrum of singular values for a model merged with task arithmetic. As the spectrum flattening can be performed on weights merged with any merging technique, we combine Iso with other merging techniques, denoting it with +Iso-C.
@@ Appendix B Computational Complexity Analysis
Iso-C Complexity. To assess the time complexity of our approach, we begin by analyzing that of Iso-C [marczak2025task], to establish a baseline for comparison. For each layer, the Iso-C procedure can be broken down into the following steps:
@@
  • SVD computation: Computing the decomposition $\Delta_{\operatorname{TA}}=U\Sigma V^{\top}$ for an $m\times n$ matrix has a complexity of $\mathcal{O}(m^{2}\cdot n+n^{3})$ [cs357_svd_2020].
@@
KnOTS Complexity. Secondly, we analyze the computational cost of the KnOTS method [stoica2024knots]:
@@
TSV Complexity. Then we analyze the cost of TSV [tsv]:
@@ C.1 High subspace alignment leads to lower interference
In this Section, we experimentally show that merging in Core Space reduces interference when merging models. We follow [yang2024representation, marczak2025task] and measure the interference as the L1 distance between the final embeddings of task-specific models and the merged one. We compare the interference when merging with TSV + Iso-C in Full Space versus Core Space. For each dataset, we collect the activations from the final layer (i.e., the projection to a common vision-language space) of both the task-specific model and the merged model. We present the average distance across all the samples in the test set. We observe lower interference when merging in Core Space, highlighting its effectiveness. Note that Full Space merging, which causes higher interference, also exhibits higher SAR in Sec.˜5.2.
@@ Appendix D Additional Experiment Details
  • KnOTS LoRA Checkpoints [stoica2024knots]: The KnOTS repository, which provides LoRA-adapted model checkpoints and training scripts, is licensed under the MIT License. This permissive license allows for reuse and modification with proper attribution.
  • Cars196 [krause20133d]: The Cars196 dataset is available for non-commercial research purposes. Specific licensing details are not explicitly provided.
  • Describable Textures Dataset (DTD) [cimpoi_describing_2014]: The DTD is made available to the computer vision community for research purposes. The dataset is licensed under the Creative Commons Attribution 4.0 License (CC BY 4.0).
  • EuroSAT [helber2018introducing]: The EuroSAT dataset is licensed under the MIT License.
  • German Traffic Sign Recognition Benchmark (GTSRB) [stallkamp2012man]: The GTSRB dataset is licensed under the Creative Commons Zero (CC0) Public Domain Dedication.
  • MNIST [lecun2010mnist]: The MNIST dataset is publicly available for research purposes. Specific licensing details are not explicitly provided; users are advised to consult the dataset’s source for more information.
  • NWPU-RESISC45 [cheng_remote_2017]: The NWPU-RESISC45 dataset is licensed under the Creative Commons Attribution 4.0 License (CC BY 4.0).
  • SUN397 [sun397]: The SUN397 dataset is available for research purposes only. Specific licensing details are not explicitly provided; users are advised to consult the dataset’s source for more information.
  • Street View House Numbers (SVHN) [svhn]: The SVHN dataset is available for non-commercial use only.
  • Stanford Natural Language Inference (SNLI) [bowman2015large]: The SNLI dataset is licensed under the Creative Commons Attribution-ShareAlike 4.0 International License (CC BY-SA 4.0).
  • Multi-Genre Natural Language Inference (MNLI) [williams2017broad1coverage]: The MNLI dataset is released under the Open American National Corpus (OANC) license, which permits free use, modification, and sharing under permissive terms.
  • Sentences Involving Compositional Knowledge (SICK) [marelli2014sick]: The SICK dataset is distributed under a Creative Commons Attribution-NonCommercial-ShareAlike license.
  • Question Natural Language Inference (QNLI) [wang2018glue]: The QNLI dataset is part of the GLUE benchmark. Specific licensing details are not explicitly provided; users are advised to consult the dataset’s source for more information.
  • Recognizing Textual Entailment (RTE) [wang2018glue]: The RTE dataset is part of the GLUE benchmark. Specific licensing details are not explicitly provided; users are advised to consult the dataset’s source for more information.
  • SciTail [scitail]: The SciTail dataset is licensed under the Apache License 2.0.
@@ D.1 Experimental Environment
Our implementation builds directly on the KnOTS codebase [stoica2024knots] and uses the exact LoRA checkpoints they released. For full details on the original training and adaptation procedures, please refer to [stoica2024knots].
@@ D.2 Hyperparameter Search
To find optimal hyperparameters for each model, we adopt the widely used validation holdout strategy [stoica2024knots, marczak2025task, tsv, yadav2023tiesmerging]. Specifically, we perform a linear search for hyperparameters on the validation set, starting from a defined minimum value and incrementally increasing it until performance declines, indicating the optimal range. The identified optimal hyperparameters are then applied to the test set. We use the following search settings:
@@
  • For CART, the pruning rank is searched over the set $\{0.04,0.08,0.16,0.32\}$, following the methodology of the original paper. Additionally, CART includes an extra scaling factor $\lambda$ in its merging formulation. Specifically, the merged weights are computed as $W_{\text{merged}}=W_{0}+\alpha(\theta_{\text{avg}}+\lambda\sum_{t=1}^{T}\bar{\tau}_{t})$, where $\theta_{\text{avg}}$ denotes the average of the updates and $\bar{\tau}_{t}$ represents the centered task vector for task $t$. For further details, we refer the reader to [choi2024revisiting].
@@ E.3 Extension to VeRA
We also evaluated the applicability of Core Space merging beyond LoRA, specifically on VeRA [kopiczko2023vera] (Tab.˜10). In VeRA, the decomposition $\Delta W=\Lambda_{b}B\Lambda_{d}A$ differs structurally from LoRA since $A$ and $B$ are fixed random matrices, and only the scaling vectors $\Lambda_{b},\Lambda_{d}$ are trainable.